# Asymptotic evaluation of the information processing capacity in reservoir computing

### v1

@@ 4 Numerical simulation
First, we show the effectiveness of our method in a system where each term of the asymptotic forms except for $o(1/T)$ can be precisely calculated. Next, we estimate the true IPCs in systems where the target outputs are given by Legendre polynomials. Finally, we apply our method to the estimation of the true IPC for the NARMA10 task.
@@ 4.1 A simple model
$x(U^{\prime})=\sum_{s=0}^{\infty}2^{-s}\,U_{-s},\ \ \hat{y}(U^{\prime})=1+x(U^{\prime}).$ (43)
The mean squared error is
@@
Thus, the true parameter is $w_{0}=1$, and the minimum value of the mean squared error is $l(w_{0})=1$. The true ICP is
@@
$C_{l,\zeta}=0,\ \ V_{l,\infty}=0,\ \ V_{\zeta,\infty}=\frac{6992}{1215},\ \ I_{\infty}=\frac{4}{3},\ \ J=\frac{4}{9}.$ (46)
We performed a numerical simulation. Training and test data were generated $10000$ times for each data length $T$, and the training and the test IPC were calculated. Their means and variances were modeled as
$C(T)=a+\frac{b}{T},\ \ C^{\prime}(T,T^{\prime})=a+\frac{c}{T^{\prime}},\ \ V(T)=\frac{d}{T},\ \ V^{\prime}(T,T^{\prime})=\frac{d}{T^{\prime}},$ (47)
and $a,b,c,d$ were estimated using the least squares method. The exact and estimated values of the asymptotic expansion are shown in Table 1.
@@
estimation 0.3076476776827442 0.1685955855549023 5.9262712596285585 1.3418518658017338
We find that the least-squares estimation is successful. We plot the means and variances of the IPC in Fig. 1 and find that they are almost on the asymptote up to the term $1/T$. Next, we plotted on the log-log scale to verify the dominant $T$-dependence. To remove the constant term, we subtracted $C_{0}$ from the means of the IPCs. (After subtracting $C_{0}$, the 18th item in the mean of the training IPC was negative, and we removed it from the log-log scale.) Most of the samples are on the asymptote up to the term $1/T$. Although the means of the training IPC fluctuate more, this is likely due to a lack of samples.
@@ 4.2 Legendre polynomials
The RC system approximates a mapping from the input to output time series by the linear combination of nonlinear dynamics. To capture the set of the functions that RC can produce, it is useful to evaluate the IPCs of the orthogonal polynomials [10, 11]. In this subsection, the input sequence $U^{\prime}$ is sampled independently from ${\rm Uniform}(-1,1)$, and the target output is the product of Legendre polynomials,
@@
We employed two target outputs. One was a short-term linear task $\hat{y}_{t}=L_{1}(u_{t-1})$, which is easy to approximate with the ESN, and the other was a long-term nonlinear task $\hat{y}_{t}=L_{15}(u_{t-5})$, which is difficult to approximate. The number of hidden nodes was $d_{1}=100$, and the ESN parameters, $v_{1},v_{2},c$, were randomly chosen from the normal distribution. The spectral radius of $v_{1}$ was set to $0.9$, and the proportion of nonzero elements was set to $0.7$. For each $T$, we generated data $1000$ times and calculated the training and the test IPC. Then, we estimated the asymptotic parameters. Figs. 2 and 3 show the results when the first- and 15th-order Legendre polynomials were used for the target outputs, respectively. The means and the variances of the true IPC for the first-order polynomial task are roughly on the asymptotic curve, indicating that the estimation was successful. In the 15th-order polynomial task, the asymptote fitted the mean IPC samples well, and the estimated true IPC was almost 00. The variances were significantly off the asymptote in the log-log graph, which reflects our mention in Section 3 that when the true IPC is 00, the variance approaches 00 faster than $1/T$.
@@
Figure 3: The means and the variances of the IPC using uniformly distributed input and 15th-order Legendre output are plotted, along with the asymptote estimated from them. Although the means are roughly on the theoretical line, the variances are significantly off. Since the true IPC is nearly 00, we can find that the variance decays faster than $1/T$. (In this graph, it decays at approximately $1/T^{2}$.)
@@ 4.3 NARMA10
The NARMA10 task is widely used as a benchmark for RC [16]. The input sequence is sampled independently from the uniform distribution. The output sequence of the NARMA10 is given by
@@
We used ${\rm Uniform}(0,0.2)$ as the input distribution and $(\alpha,\beta,\gamma,\delta)=(0.3,0.05,1.5,0.1)$ as NARMA10 parameters. The number of hidden nodes was $d_{1}=100$, and the ESN parameters $v_{1},v_{2},c$, were randomly chosen from the normal distribution. The spectral radius of $v_{1}$ was set to $0.9$, and the proportion of non-zero elements was set to $0.7$. For each $T$, data were generated $1000$ times, and the training and the test IPC were calculated. Then, we estimate the asymptotic parameters. The results are shown in Fig. 4. We can see that both the means and variances are almost on the asymptote.
@@ 5 Summary
RNNs and their variants such as LSTM [4] and GRU [5] can approximate the relationship between the input and the output time series. However, the gradient learning takes a long time as data length becomes longer. RC is a system that approximates the relationship between input and output time series by the linear combination of nonlinear dynamics and finishes training quickly by optimizing only the linear combination.
The performance of RC is evaluated by the mean squared error or the squared error normalized by the target output. The latter is called the IPC [10, 11]. To know the approximation accuracy of the relationship between input and output time series, we should evaluate the IPC for infinitely long data, not the IPC for finite-length data. Simply evaluating the IPC with long data cannot remove the difference from the limit value. We propose a method to estimate the IPC for infinitely long data by using the asymptotic expansions of the training and the test IPC and the least-squares method. Then, we show the validity of our method by numerical simulations. Although we used ESNs in the demonstrations, our method can be applied to RC systems that satisfies the finiteness of the covariance matrices, as mentioned in the last second paragraph of Section 3.
Our method can estimate the IPC more accurately than simply extending the data length. However, it is computationally heavy due to multiple trials with long data. This is a trade-off with the estimation accuracy. A computationally less costly method for evaluating the IPC is our future work.
@@ Acknowledgement
This paper is based on results obtained from a project, JPNP16007, commissioned by the New Energy and Industrial Technology Development Organization (NEDO). We appreciate Professor Takashi Morie and Doctor Tomoyuki Kubota for their fruitful comments.
Appendix A Least mean square
Let $f_{1}(n)=a+\frac{b_{1}}{n}$ and $f_{2}(n)=a-\frac{b_{2}}{n}$ be the asymptotic mean IPCs for the training and the test data, respectively, and $\{(n_{i,1},g_{i,1})\}_{i=1}^{N}$ and $\{(n_{i,2},g_{i,2})\}_{i=1}^{N}$ are the mean IPCs for the training and the test data at each data length. We determine $a,b_{1},b_{2}$ by the least mean square of the cost function below,
$L(a,b,b^{\prime})=\frac{1}{2}\sum_{i=1}^{N}\left[\left(a+\frac{b_{1}}{n_{i,1}}-g_{i,1}\right)^{2}+\left(a-\frac{b_{2}}{n_{i,2}}-g_{i,2}\right)^{2}\right].$ (50)
Using the notation,
$\alpha_{1}=\sum_{i=1}^{N}\frac{1}{n_{i,1}},\ \ \alpha_{2}=\sum_{i=1}^{N}\frac{1}{n_{i,2}},\ \ \beta_{1}=\sum_{i=1}^{N}\frac{1}{n_{i,1}^{2}},\ \ \beta_{2}=\sum_{i=1}^{N}\frac{1}{n_{i,2}^{2}},$ (51)
$s_{1}=\sum_{i=1}^{N}g_{i,1},\ \ s_{2}=\sum_{i=1}^{N}g_{i,2},\ \ t_{1}=\sum_{i=1}^{N}\frac{g_{i,1}}{n_{i,1}},\ \ t_{2}=\sum_{i=1}^{N}\frac{g_{i,2}}{n_{i,2}},$ (52)
we can obtain $a,b,b^{\prime}$ from
$\begin{pmatrix}2N&\alpha_{1}&-\alpha_{2}\\ \alpha_{1}&\beta_{1}&0\\ \alpha_{2}&0&-\beta_{2}\\ \end{pmatrix}\,\begin{pmatrix}a\\ b_{1}\\ b_{2}\\ \end{pmatrix}=\begin{pmatrix}s_{1}+s_{2}\\ t_{1}\\ t_{2}\\ \end{pmatrix}.$ (53)
@@ Appendix A Least mean square
In the same way, let $f^{\prime}_{1}(n)=\frac{d}{n}$ and $f^{\prime}_{2}(n)=\frac{d}{n}$ are the asymptotic IPC variances for the training and the test data, respectively, and $\{(n_{i,1},g^{\prime}_{i,1})\}_{i=1}^{N}$ and $\{(n_{i,2},g^{\prime}_{i,2})\}_{i=1}^{N}$ are the IPC variances at each data length. We determine $d$ by the least mean square,
$L^{\prime}(d)=\frac{1}{2}\sum_{i=1}^{N}\left[\left(\frac{d}{n_{i,1}}-g^{\prime}_{i,1}\right)^{2}+\left(\frac{d}{n_{i,2}}-g^{\prime}_{i,2}\right)^{2}\right],$ (54)
@@
$d=\frac{\sum_{i=1}^{N}\left(\frac{g^{\prime}_{i,1}}{n_{i,1}}+\frac{g^{\prime}_{i,2}}{n_{i,2}}\right)}{\sum_{i=1}^{N}\left(\frac{1}{n_{i,1}^{2}}+\frac{1}{n_{i,2}^{2}}\right)}=\frac{t^{\prime}_{1}+t^{\prime}_{2}}{\beta_{1}+\beta_{2}},$ (55)
where
$t^{\prime}_{1}=\sum_{i=1}^{N}\frac{g^{\prime}_{i,1}}{n_{i,1}},\ \ t^{\prime}_{2}=\sum_{i=1}^{N}\frac{g^{\prime}_{i,2}}{n_{i,2}}.$ (56)

### v2

@@ 4 Numerical simulation
First, we show the effectiveness of our method in a system where each term of the asymptotic forms except for $o(1/T)$ can be precisely calculated. Next, we estimate the true IPCs in systems where the target outputs are given by Legendre polynomials. Finally, we apply our method to the estimation of the true IPC for the NARMA10 task. In these simulations, we compared the proposed method with the method in Ref. [dambre2012information]. The threshold in Eq. (8) is given by $\chi^{2}$ distribution as
$C_{T,{\rm th}}=\frac{2\alpha}{T},\ \ {\rm Prob}(\chi^{2}(N)\geq\alpha)=p,$ (55)
where $p$ is a small positive real number, and we set $p=10^{-4}$.
@@ 4.1 A simple model
$x(u^{\prime}_{t})=\sum_{s=0}^{\infty}2^{-s}\,u_{t-s},\ \ \hat{y}(u^{\prime}_{t})=1+x(u^{\prime}_{t}).$ (56)
The MSE is
@@
Thus, the true parameter is $w_{0}=1$, and the minimum value of the MSE is $l(w_{0})=1$. The true IPC is
@@
$C_{l,\mu}=0,\ \ V_{l,\infty}=0,\ \ V_{\mu,\infty}=\frac{6992}{1215},\ \ I_{\infty}=\frac{4}{3},\ \ J=\frac{4}{9}.$ (59)
We performed a numerical simulation. Training and test data were generated $10^{5}$ times for each data length $T$, and the training and the test IPC were calculated. Their means and variances were modeled as
$C(T)=a+\frac{b_{1}}{T},\ \ C^{\prime}(T,T^{\prime})=a-\frac{b_{2}}{T^{\prime}},\ \ V(T)=\frac{d}{T},\ \ V^{\prime}(T,T^{\prime})=\frac{d}{T^{\prime}},$ (60)
and $a,b_{1},b_{2},d$ were estimated using the weighted least squares method. To confirm the effectiveness of the asymptotic expansion, we used the first $5$ training and test data, that is, $T=200,\ldots,600$ and $T^{\prime}=400,\ldots,1200$. The exact and estimated values of the asymptotic expansion are shown in Table 1.
@@
estimation 0.3075 0.1933 5.8906 1.412
We find that the weighted least-squares estimation is successful. We plot the means and variances of the IPC in Fig. 1 and find that they are almost on the asymptote up to the term $1/T$. Next, we plotted on the log-log scale to verify the dominant $T$-dependence. To remove the constant term, we subtracted $C_{0}$ from the means of the IPCs. (After subtracting $C_{0}$, the 17th item in the mean of the training IPC was negative, and we removed it from the log-log scale.) Most of the samples are on the asymptote up to the term $1/T$, which shows the validity of the asymptotic expansion. Although the means of the training IPC fluctuate more, this is likely due to a lack of samples.
Finally, we compare the proposed method with the method in Ref. [dambre2012information]. We used the sample mean of the training IPC at $T=600$. Since the threshold is $C_{T,{\rm th}}=0.0505$, the empirical IPC is $0.3080$. Our result is slightly closer to the ground truth, showing that the proposed method is effective (the differences from the ground truth are $2\times 10^{-4}$ and $3\times 10^{-4}$, respectively).
@@ 4.2 Legendre polynomials
An RC system approximates a mapping from input to output time series by the linear combination of nonlinear dynamics. To capture the set of functions that RC can produce, it is useful to evaluate the IPCs of the orthogonal polynomials [dambre2012information, kubota2021unifying]. In this subsection, the input sequence $U^{\prime}$ is sampled independently from ${\rm Uniform}(-1,1)$, and the target output is the product of Legendre polynomials,
@@
We employed two target outputs. One was a short-term linear task $\hat{y}_{t}=L_{1}(u_{t-1})$, which is easy to approximate with the ESN, and the other was a long-term nonlinear task $\hat{y}_{t}=L_{15}(u_{t-5})$, which is difficult to approximate. The number of hidden nodes was $d_{1}=100$, the ESN parameters, $v_{1},v_{2}$, were randomly chosen from the normal distribution, and $c=0$. The spectral radius of $v_{1}$ was set to $0.9$, and the proportion of nonzero elements was set to $0.7$. For each $T$, we generated data $10^{3}$ times and calculated the training and the test IPC. To confirm the usefulness of the asymptotic expansion, we used the training and test data from $T=10^{3}$ to $T=10^{4}$ and estimated the asymptotic parameters. We obtained $a=0.997632,b_{1}=0.242030,b_{2}=0.265197,d=1.6\times 10^{-5}$ for the first-order task and $a=0.00070,b_{1}=99.63334,b_{2}=110.70010,d=0.13467$ for the 15th-order task. Figs. 2 and 3 show the results when the first- and 15th-order Legendre polynomials were used for the target output, respectively. The means and the variances of the true IPC for the first-order polynomial task are roughly on the asymptotic curve, indicating that the estimation was successful. In the 15th-order polynomial task, the asymptote fitted the mean IPC samples well, and the estimated true IPC was almost 0. The variances were significantly off the asymptote in the log-log graph, which reflects our mention in the last paragraph of subsection 3.1 that when the true IPC is 0, the variance approaches 0 faster than $1/T$.
Finally, we compare our method with the method in Ref. [dambre2012information]. We used the sample mean of the training IPC at $T=10^{4}$. Since the threshold is $C_{T,{\rm th}}=0.032264$, the empirical IPC for the target output $\hat{y}_{t}=L_{1}(u_{t-1})$ is $0.997654$. For the target output $\hat{y}_{t}=L_{15}(u_{t-1})$, the sample mean of the training IPC is $0.01017$, which is below the threshold, and therefore the empirical IPC is 0. Since the ground truth is unknown, it is difficult to compare the results. In the first-order task, the proposed method is closer to the IPC $0.997631$ estimated by longer data, $T=10^{5}$, which shows the usefulness of the asymptotic expansion. In the 15th-order task, both methods produced the same IPC. When determining whether the IPC is 0, it is considered more robust to use the power of the variance, which is less susceptible to fluctuations caused by finite-length data than the numerical value of the IPC.
@@
Figure 3: The means and the variances of the IPC using uniformly distributed input and 15th-order Legendre output are plotted, along with the asymptote estimated from them. Although the means are roughly on the theoretical line, the variances are significantly off. Since the true IPC is nearly 0, we can find that the variance decays faster than $1/T$. (In this graph, it decays at approximately $1/T^{2}$.)
@@ 4.3 NARMA10
The NARMA10 task is widely used as a benchmark for RC [atiya2000new]. The input sequence is sampled independently from the uniform distribution. The output sequence of the NARMA10 is given by
@@
We used ${\rm Uniform}(0,0.2)$ as the input distribution and $(\alpha,\beta,\gamma,\delta)=(0.3,0.05,1.5,0.1)$ as NARMA10 parameters. The number of hidden nodes was $d_{1}=100$, the ESN parameters $v_{1},v_{2}$, were randomly chosen from the normal distribution, and $c=0$. The spectral radius of $v_{1}$ was set to $0.9$, and the proportion of nonzero elements was set to $0.7$. For each $T$, we generated the data $10^{3}$ times and calculated the training and the test IPC. To confirm the usefulness of the asymptotic expansion, we used the training and test data from $T=10^{3}$ to $T=10^{4}$ and estimated the asymptotic parameters. We obtained $a=0.998283,b_{1}=0.193536,b_{2}=0.215439,d=2.0\times 10^{-5}$. The results are shown in Fig. 4. We can find that both the means and variances are almost on the asymptote.
Finally, we compare our method with the method in Ref. [dambre2012information]. We used the sample mean of the training IPC at $T=10^{4}$. Since the threshold is $C_{T,{\rm th}}=0.032264$, the empirical IPC is $0.9983023$. Due to the unknown ground truth, it is difficult to compare the results. However, the proposed method is closer to the IPC $0.998281$ estimated by longer data, $T=10^{5}$, which shows the usefulness of the asymptotic expansion.
@@ 5 Summary
RNNs and their variants such as LSTM [hochreiter1997long] and GRU [cho2014learning] can approximate relationship between input and output time series. However, gradient learning takes a long time as the data length is longer. RC is a system that approximates the relationship between input and output time series by the linear combination of nonlinear dynamics and finishes training quickly by optimizing only the linear combination.
The performance of RC is evaluated by the MSE or the squared error normalized by the target output. The latter is called the IPC [dambre2012information, kubota2021unifying]. To know the approximation accuracy of the relationship between input and output time series, we should evaluate the IPC for infinitely long data, not the IPC for finite-length data. Simply evaluating the IPC with long data cannot remove the difference from the limit value. We proposed a method to estimate the IPC for infinitely long data by using the asymptotic expansions of the training and the test IPC and the weighted least-squares method. We also explained that to determine whether the IPC is 0, the asymptotic expansion of the variance of the IPC is a robust method. Then, we showed the validity of our method by numerical simulations. Although we used ESNs in the demonstrations, our method can be applied to RC systems that satisfy the conditions mentioned in the last paragraph of Section 3. Since the numerator of the IPC contains the MSE, the proposed method can be applied to estimate the MSE the infinite time limit.
Our method can estimate the IPC more accurately than simply using long data. The drawback of the proposed method is that it is computationally intensive due to multiple trials on long data. If we naively implement the proposed method, the computation complexity is approximately $10^{3}$ times $O(NTd_{1}^{3})$. Although we have explained a strategy for reducing the amount of calculation in subsection 3.2, we cannot expect a drastic reduction in computation time, and therefore, the proposed method is considered applicable up to about $d_{1}\sim 10^{3}$.
We consider a more practical online estimation method which uses Eq. (40) without taking the sample means. Since it is computationally expensive to perform inverse matrix calculations for each $T$ to find $w_{T}$, we substitute it with $\widetilde{w}_{T}$ obtained by the gradient method with a stochastic approximation [robbins1951stochastic, yuan2024optoelectronic]. In Eq. (40), $\bar{l}_{T}(w),\bar{\mu}_{T}$, and $\xi_{T}$ follow a normal distribution for a sufficiently large $T$, and $\widetilde{w}_{T}$ converges to the true parameter $w_{0}$. Thus, if we can estimate the mean of the normal distributions using these terms at different $T$, we can also estimate the true IPC. We will leave these for future work.
@@ Acknowledgement
This paper is based on results obtained from a project, JPNP16007, commissioned by the New Energy and Industrial Technology Development Organization (NEDO). We appreciate Professor Takashi Morie, Doctor Tomoyuki Kubota, and Doctor So Nakashima for their fruitful comments.
Appendix A Weighted least squares method
First, we derive the asymptotic form of the sample mean of the training IPC. Suppose that the number of samples $N$ is large enough. Applying the CLT to the sample mean of each term in Eq. (41), we obtain
$f_{1}(T)=$ $1-\frac{l(w_{0})}{\mu_{0}}+\frac{1}{T}\left[\frac{C_{l,\mu}}{\mu_{0}^{2}}-\frac{l(w_{0})\,V_{\mu,\infty}}{\mu_{0}^{3}}+\frac{{\rm Tr}(I_{\infty}J^{-1})}{\mu_{0}}\right]+\epsilon_{T}+o_{p}\left(\frac{1}{T}\right)$
$=$ $a+\frac{b_{1}}{T}+\epsilon_{T}+o_{p}\left(\frac{1}{T}\right),$ (63)
$\epsilon_{T}\sim$ ${\cal N}\left(0,\frac{1}{NT}\left[\frac{V_{l,\infty}}{\mu_{0}^{2}}+\frac{l(w_{0})^{2}\,V_{\mu,\infty}}{\mu_{0}^{4}}-\frac{2}{\mu_{0}^{3}}\,l(w_{0})\,C_{l,\mu}\right]\right).$ (64)
Similarly, the asymptotic form of the sample mean of the test IPC can be written as
$f_{2}(T)=$ $1-\frac{l(w_{0})}{\mu_{0}}+\frac{1}{T}\left[\frac{T}{T^{\prime}}\,\frac{C_{l,\mu}}{\mu_{0}^{2}}-\frac{T}{T^{\prime}}\,\frac{l(w_{0})\,V_{\mu,\infty}}{\mu_{0}^{3}}-\frac{{\rm Tr}(I_{\infty}J^{-1})}{\mu_{0}}\right]+\epsilon^{\prime}_{T^{\prime}}+o_{p}\left(\frac{1}{T}\right)$
$=$ $a+\frac{b_{2}}{T}+\epsilon^{\prime}_{T^{\prime}}+o_{p}\left(\frac{1}{T}\right),$ (65)
$\epsilon^{\prime}_{T^{\prime}}\sim$ ${\cal N}\left(0,\frac{1}{NT}\left[\frac{T}{T^{\prime}}\,\frac{V_{l,\infty}}{\mu_{0}^{2}}+\frac{T}{T^{\prime}}\,\frac{l(w_{0})^{2}\,V_{\mu,\infty}}{\mu_{0}^{4}}-\frac{T}{T^{\prime}}\,\frac{2}{\mu_{0}^{3}}\,l(w_{0})\,C_{l,\mu}\right]\right).$ (66)
Next, we obtain $a,b_{1}$, and $b_{2}$ using the weighted least squares method. Let $\{(T_{i,1},g_{i,1})\}_{i=1}^{N}$ and $\{(T_{i,2},g_{i,2})\}_{i=1}^{N}$ be the mean IPCs for the training and the test data for each data length. We assume that all data lengths are sufficiently long so that the $o_{p}(1/T)$ terms in Eqs. (63) and (65) can be neglected. The least squares method is the maximum likelihood estimation when a normal distribution is used for the likelihood function. Since the variance of $\epsilon_{T}$ and $\epsilon^{\prime}_{T^{\prime}}$ is proportional to $1/T$, we use the following cost function, which cancels the change in variance due to the data length;
$L(a,b,b^{\prime})=\frac{1}{2}\sum_{i=1}^{N}\left[T_{i,1}\left(a+\frac{b_{1}}{T_{i,1}}-g_{i,1}\right)^{2}+T_{i,2}\left(a-\frac{b_{2}}{T_{i,2}}-g_{i,2}\right)^{2}\right].$ (67)
Eq. (67) can be interpreted as a cost function that places more emphasis on fitting to the data with less fluctuation. Using the notation,
$\beta_{1}=\sum_{i=1}^{N}\frac{1}{T_{i,1}},\ \ \beta_{2}=\sum_{i=1}^{N}\frac{1}{T_{i,2}},\ \ s_{1}=\sum_{i=1}^{N}T_{i,1}g_{i,1},\ \ s_{2}=\sum_{i=1}^{N}T_{i,2}g_{i,2},$ (68)
$t_{1}=\sum_{i=1}^{N}g_{i,1},\ \ t_{2}=\sum_{i=1}^{N}g_{i,2},\ \ \gamma=\sum_{i=1}^{N}\left(T_{i,1}+T_{i,2}\right),$ (69)
we can obtain $a,b_{1},b_{2}$ from
$\begin{pmatrix}\gamma&N&-N\\ N&\beta_{1}&0\\ N&0&-\beta_{2}\\ \end{pmatrix}\,\begin{pmatrix}a\\ b_{1}\\ b_{2}\\ \end{pmatrix}=\begin{pmatrix}s_{1}+s_{2}\\ t_{1}\\ t_{2}\\ \end{pmatrix}$. (70)
Next, we evaluate the variances of the training IPC and the test IPC. First, we estimate the fluctuation in the unbiased variance of the training IPC. Applying the CLT to Eq. (40), we find that
$\sqrt{T}(C_{T}(U^{\prime}_{T})-\mathbb{E}[C_{T}(U^{\prime}_{T})])\xrightarrow{d}{\cal N}(0,\alpha),$ (71)
$\alpha=\frac{V_{l,\infty}}{\mu_{0}^{2}}+\frac{l(w_{0})^{2}V_{\mu,\infty}}{\mu_{0}^{4}}-\frac{2}{\mu_{0}^{3}}l(w_{0})C_{l,\mu}.$ (72)
The sample mean of $(T/\alpha)\,(C_{T}(U^{\prime}_{T})-\mathbb{E}[C_{T}(U^{\prime}_{T})])^{2}$ follows $\chi^{2}(N)$. Thus, the distribution of the unbiased variance of $C_{T}$ can be approximated by the normal distribution whose variance is $(2N\alpha^{2})/((N-1)^{2}T^{2})$.
@@ Appendix A Weighted least squares method
Let $f^{\prime}_{1}(T)=\frac{d}{T}$ and $f^{\prime}_{2}(T)=\frac{d}{T}$ be the asymptotic IPC variances for the training and the test data, respectively, and $\{(T_{i,1},\sigma^{2}_{i,1})\}_{i=1}^{N}$ and $\{(T_{i,2},\sigma^{2}_{i,2})\}_{i=1}^{N}$ be the IPC variances for each data length. We employ the cost function with the weight of the variance,
$L^{\prime}(d)=\frac{1}{2}\sum_{i=1}^{N}\left[T_{i,1}^{2}\left(\frac{d}{T_{i,1}}-\sigma^{2}_{i,1}\right)^{2}+T_{i,2}^{2}\left(\frac{d}{T_{i,2}}-\sigma^{2}_{i,2}\right)^{2}\right],$ (73)
@@
$d=\frac{\sum_{i=1}^{N}(T_{i,1}\sigma^{2}_{i,1}+T_{i,2}\sigma^{2}_{i,2})}{2N}.$ (74)